calculate the number of moles of oxygen required to to yield 72g of water


1
Expert's answer
2021-03-09T06:06:44-0500

Mass of Water = 72 g

Molar mass of water(H2O) = 18 g/mol

Moles of water produced = "\dfrac{72}{18}=4\space mol"

"H_2+\dfrac{1}{2}O_2\longrightarrow H_2O"

According to balanced chemical reaction,

1 mol H2O is produced by the reaction of 1/2 mol O2

4 mol H2O is produced by the reaction of "4\times\dfrac{1}{2}=2\space mol" O2

Moles of O2 required = 2 moles